# Refined Bitcoin Security-Latency Under Network Delay

### v3

@@ 2.3 Parameters and Definitions
(Block publication) We say a block becomes public when it enters the view of every honest miner. Any honest block mined at $\tau_{m}$, becomes public by $\tau_{m}+\Delta$, i.e., enters the view of all honest nodes by $\tau_{m}+\Delta$.
@@
$\rho$ $\alpha e^{-\lambda\Delta}(1+\lambda\Delta+\beta-\alpha e^{-\lambda\Delta})$
@@
$\bar{a}_{i}$ $\sum_{j\geq i}\bar{\alpha}_{j}+\beta*\mathbbm{1}_{i\leq 2}$
$\bar{b}_{i}$ $\sum_{j\geq i}\bar{\alpha}_{j}+\beta*\mathbbm{1}_{i\leq 1}$
$a_{i}$ $\sum_{j\geq i}\alpha_{j}+\beta*\mathbbm{1}_{i\leq 2}$
$b_{i}$ $\sum_{j\geq i}\alpha_{j}+\beta*\mathbbm{1}_{i\leq 1}$
@@ 3.1 Pre-Mining Gain
Letting $\pi_{i}$ denote $\mathbb{P}(L=i)$, we find $\pi_{0}=\frac{\alpha_{0}-\beta_{1}}{1-\beta_{1}}$, $\pi_{1}=\pi_{0}\cdot\frac{1-\alpha_{0}}{\alpha_{0}}$, and $\pi_{i}=\pi_{1}(\frac{\beta_{1}}{\alpha_{0}})^{i-1}$ for $i>1$. We denote the PMF of $L$ with $P_{1}$, CDF of $L$ with $F_{1}$, and CCDF of $L$ with $\bar{F}_{1}=1-F_{1}$.
@@ 3.3 Post-Confirmation Race
Consider a binary random walk starting from the origin and moving to the left w.p. $\alpha$ and right w.p. $\beta$. Let $M$ denote the maximum point the random walk ever reaches. Let $N$ denote the number of times the random walk reaches the point $M$. Then,
@@
Consider a random walk starting from the origin with a three-way coin toss, each toss denoted by $W_{i}$, $i\geq 1$, and $\mathbb{P}(W_{i}=-1)=\mathbb{P}(E_{1})=\alpha_{0}$, $\mathbb{P}(W_{i}=0)=\mathbb{P}(E_{2})=\alpha_{1}$, $\mathbb{P}(W_{i}=1)=\mathbb{P}(E_{3})=\beta_{1}$. Let $M$ denote the maximum point the random walk ever reaches. Let $N$ denote the number of times the random walk moves from $M-1$ to $M$. Then,
@@
Consider a sequence of i.i.d. random variables denoted by $W_{i}$, $i\geq 1$, and $\mathbb{P}(W_{i}=-1)=\mathbb{P}(E_{1})=\alpha_{0}$, $\mathbb{P}(W_{i}=0)=\mathbb{P}(E_{2})=\alpha_{1}$, $\mathbb{P}(W_{i}=1)=\mathbb{P}(E_{3})=\beta_{1}$. Let $T^{\prime}_{i}=\mathds{1}_{W_{i}=0}+\sum_{j=1}^{i}W_{j}$. Then, for $a\geq 1$,
@@
As mentioned previously, $\max\limits_{i\geq 1}T^{\prime}_{i}\geq k-L-S$ refers to the event that the adversary catches the honest chain during the post-confirmation race. We denote the PMF of $\max\limits_{i\geq 1}T^{\prime}_{i}$ with $P_{3}$. Finally, we put everything together, namely, the analysis of the steady state (i.e., pre-mining gain), confirmation interval, and post-confirmation race in the following theorem.
@@
$\bar{F}_{1}(k\!-\!1)+\sum_{i<k}P_{1}(i)\bar{F}_{2}(k\!-\!1\!-\!i)+\sum_{i+j<k}P_{1}(i)P_{2}(j)\bar{F}_{3}(k\!-\!1\!-\!i\!-\!j)$ (33)
@@ 4.1 Pre-Mining Gain
Steady state of $\bar{P}$ can be found recursively using
@@
where $\pi_{0}=\frac{1-2\beta-\alpha\lambda\Delta}{\alpha}$.
@@
We denote the upper bound on the lead as $\bar{L}$ and the PMF of $\bar{L}$ with $P^{\prime}_{1}$.
@@ 4.3 Post-Confirmation Race
where $\bar{M}$ denotes the maximum point this random walk ever reaches. Let $\bar{M}^{\prime}$ denote the maximum difference between the heights of adversarial and honest chains under the simplified $(A,H)$ random walk model with two arrivals at a time. Thus, if $\bar{D}$ is even, then we have $\mathbb{P}(\bar{M}\geq\bar{D})=\mathbb{P}(\bar{M}^{\prime}\geq\bar{D})$. If odd, however, there can be cases of unobservable adversarial win at the end of two tosses. To avoid this complication, we simply consider a single toss (using the rigged model of [5]) initially to make sure deficit becomes even before using (58): $\mathbb{P}(\bar{M}^{\prime}\geq\bar{D})=\bar{\alpha}_{0}\mathbb{P}(\bar{M}\geq\bar{D}+1)+(1-\bar{\alpha}_{0})\mathbb{P}(\bar{M}\geq\bar{D}-1)$. We denote PMF of $\bar{M}^{\prime}$ with $P^{\prime}_{3}$.
@@
$\bar{F}^{\prime}_{1}(k\!-\!1)+\sum_{i<k}P^{\prime}_{1}(i)\bar{F}^{\prime}_{2}(k\!-\!1\!-\!i)+\sum_{i+j<k}P^{\prime}_{1}(i)P^{\prime}_{2}(j)\bar{F}^{\prime}_{3}(k\!-\!1\!-\!i\!-\!j)$ (59)
@@ Appendix A Appendix
(Ramaswami[8]) Steady state of $P$ can be found recursively using
@@
Note that with the random walk presented in $\bar{P}$, the random walk escapes to infinity when $1\leq 2\beta+\alpha\lambda\Delta$. Hence, the maximum provable fault tolerance for the rigged model using $\bar{P}$ is $1>2\beta+\alpha\lambda\Delta$, whereas the ultimate fault tolerance is $1>2\beta+\alpha\beta\lambda\Delta$. However, as we use the simple rigged model partially in the post-confirmation race when we find the upper bound, $\frac{1}{2}<\alpha e^{-\lambda\Delta}$ is the fault tolerance of Theorem 2.
Further, with the random walk presented in $P$, the random walk escapes to infinity when $1\leq\beta(2+\alpha\lambda\Delta)$ which is the same condition as $\beta\geq\frac{1-\beta}{1+(1-\beta)\lambda\Delta}$, i.e., the converse of the ultimate fault tolerance condition. This is not surprising, since $P$ presents the $\Delta$ delay strategy, where the honest chain grows at the slowest possible rate, i.e., $\frac{1}{\Delta+1/\alpha\lambda}$, and is in a race with adversarial chain that has the rate $\beta\lambda$.
In our figures, we use the steady states of $P^{\prime}$ and $\bar{P}$ consistently. We note that the bounds of the two lower bounds visually overlap with each other, hence we refrain from comparing them with plots. As an example, for Bitcoin $6$-block confirmation rule with $\alpha=0.9$, using $P^{\prime}$ results in safety violation lower bound probability of $0.00112412$ whereas $P$ gives $0.00112415$.

### v4

@@ 2.3 Parameters and Definitions
(Block publication) We say a block becomes public when it enters the view of every honest miner. Any honest block mined at $t$, becomes public by $t+\Delta$, i.e., enters the view of all honest nodes by $t+\Delta$.
@@
$\rho$ $\alpha e^{-\lambda\Delta}\cdot(1+\lambda\Delta+\beta-\alpha e^{-\lambda\Delta})$
@@
$\bar{a}_{i}$ $\sum_{j\geq i}\bar{\alpha}_{j}+\beta\cdot\mathbbm{1}_{i\leq 2}$
$\bar{b}_{i}$ $\sum_{j\geq i}\bar{\alpha}_{j}+\beta\cdot\mathbbm{1}_{i\leq 1}$
$a_{i}$ $\sum_{j\geq i}\alpha_{j}+\beta\cdot\mathbbm{1}_{i\leq 2}$
$b_{i}$ $\sum_{j\geq i}\alpha_{j}+\beta\cdot\mathbbm{1}_{i\leq 1}$
@@ 3.1 Pre-Mining Gain
Letting $\pi_{i}$ denote $\mathbb{P}(L=i)$, we find $\pi_{0}=\frac{\alpha_{0}-\beta_{1}}{1-\beta_{1}}$, $\pi_{1}=\pi_{0}\cdot\frac{1-\alpha_{0}}{\alpha_{0}}$, and $\pi_{i}=\pi_{1}(\frac{\beta_{1}}{\alpha_{0}})^{i-1}$ for $i>1$. We denote the PMF of $L$ with $P_{1}$.
@@ 3.3 Post-Confirmation Race
Consider a binary random walk starting from the origin and moving to the left w.p. $\alpha$ and right w.p. $\beta<\frac{1}{2}$. Let $M$ denote the maximum point the random walk ever reaches. Let $N$ denote the number of times the random walk reaches the point $M$. Then,
@@
Consider a random walk starting from the origin with a three-way coin toss, each toss denoted by $W_{i}$, $i\geq 1$, and $\mathbb{P}(W_{i}=-1)=\mathbb{P}(E_{1})=\alpha_{0}$, $\mathbb{P}(W_{i}=0)=\mathbb{P}(E_{2})=\alpha_{1}$, $\mathbb{P}(W_{i}=1)=\mathbb{P}(E_{3})=\beta_{1}$, where $\beta_{1}<\alpha_{0}$. Let $M$ denote the maximum point the random walk ever reaches. Let $N$ denote the number of times the random walk moves from $M-1$ to $M$. Then,
@@
Consider a sequence of i.i.d. random variables denoted by $W_{i}$, $i\geq 1$, and $\mathbb{P}(W_{i}=-1)=\mathbb{P}(E_{1})=\alpha_{0}$, $\mathbb{P}(W_{i}=0)=\mathbb{P}(E_{2})=\alpha_{1}$, $\mathbb{P}(W_{i}=1)=\mathbb{P}(E_{3})=\beta_{1}$ where $\beta_{1}<\alpha_{0}$. Let $T^{\prime}_{i}=\mathds{1}_{W_{i}=0}+\sum_{j=1}^{i}W_{j}$. Then, for $a\geq 1$,
@@
As mentioned previously, $\max\limits_{i\geq 1}T^{\prime}_{i}\geq k-L-S$ refers to the event that the adversary catches the honest chain during the post-confirmation race. We note that PMF of $\max\limits_{i\geq 1}T^{\prime}_{i}$ and $L$ coincide, hence we denote it with $P_{1}$ as well. Finally, we put everything together, namely, the analysis of the steady state (i.e., pre-mining gain), confirmation interval, and post-confirmation race in the following theorem.
@@
$1-\sum_{i+j+l<k}P_{1}(i)P_{2}(j)P_{1}(l)$ (33)
where
$P_{1}(0)=\frac{\alpha_{0}-\beta_{1}}{1-\beta_{1}},$
$P_{1}(i)=P_{1}(0)\cdot\frac{1-\alpha_{0}}{\alpha_{0}}\cdot(\frac{\beta_{1}}{\alpha_{0}})^{i-1},\quad i\geq 1,$
$P_{2}(j)=\alpha_{0}^{k}\beta^{j}\sum_{n=0}^{j}\binom{k\!-\!1\!+\!n}{n}\frac{(\lambda\Delta k)^{j-n}}{(j-n)!},\quad j\geq 0$
where $\beta_{1}<\alpha_{0}$.
@@ 4.1 Pre-Mining Gain
For $1>2\beta+\alpha\lambda\Delta$, steady state of $\bar{P}$ can be found recursively using
@@
where $\pi_{0}=\frac{1-2\beta-\alpha\lambda\Delta}{\alpha}$, $\bar{a}_{i}=\sum_{j\geq i}\bar{\alpha}_{j}+\beta\cdot\mathbbm{1}_{i\leq 2}$ and $\bar{b}_{i}=\sum_{j\geq i}\bar{\alpha}_{j}+\beta\cdot\mathbbm{1}_{i\leq 1}$.
@@
We denote the upper bound on the lead as $\bar{L}$ and the PMF of $\bar{L}$ with $P^{\prime}_{1}$. Note that with the random walk presented in $\bar{P}$, the random walk escapes to infinity when $1\leq 2\beta+\alpha\lambda\Delta$. Hence, the maximum provable fault tolerance for the rigged model using $\bar{P}$ is $1>2\beta+\alpha\lambda\Delta$, whereas the ultimate fault tolerance is $1>2\beta+\alpha\beta\lambda\Delta$.
@@ 4.3 Post-Confirmation Race
where $\bar{M}$ denotes the maximum point this random walk ever reaches. Let $\bar{M}^{\prime}$ denote the maximum difference between the heights of adversarial and honest chains under the simplified $(A,H)$ random walk model with two arrivals at a time. Thus, if $\bar{D}$ is even, then we have $\mathbb{P}(\bar{M}\geq\bar{D})=\mathbb{P}(\bar{M}^{\prime}\geq\bar{D})$. If odd, however, there can be cases of unobservable adversarial win at the end of two tosses. To avoid this complication, we simply consider a single toss (using the rigged model of [5]) initially to make sure deficit becomes even before using (58): $\mathbb{P}(\bar{M}^{\prime}\geq\bar{D})=\bar{\alpha}_{0}\mathbb{P}(\bar{M}\geq\bar{D}+1)+(1-\bar{\alpha}_{0})\mathbb{P}(\bar{M}\geq\bar{D}-1)$. We denote CDF of $\bar{M}^{\prime}$ with $F^{\prime}_{3}$.
@@
$1-\sum_{i+j<k}P^{\prime}_{1}(i)P^{\prime}_{2}(j)F^{\prime}_{3}(k-1-i-j)$ (59)
where
$P^{\prime}_{1}(0)=\frac{1-2\beta-\alpha\lambda\Delta}{\alpha},$
$P^{\prime}_{1}(i)=\frac{P^{\prime}_{1}(0)\bar{b}_{i}+\sum_{j=1}^{i-1}P^{\prime}_{1}(j)\bar{a}_{i+1-j}}{1-\bar{a}_{1}},\quad i\geq 1,$
$P^{\prime}_{2}(j)=\bar{\alpha}_{0}^{k}\sum_{n=0}^{j}\binom{k\!-\!1\!+\!n}{n}\frac{(\lambda\Delta k)^{j-n}}{(j-n)!}\beta^{n},\quad j\geq 0,$
[…]
where $\bar{\beta}<\bar{\alpha}$.
Although the maximum provable fault tolerance for the rigged model using $\bar{P}$ is $1>2\beta+\alpha\lambda\Delta$, as we use a simpler model in the post-confirmation race when we find the upper bound where adversary has even more advantage, $\bar{\beta}<\bar{\alpha}$ is the fault tolerance of Theorem 2.
5 Numerical Results
@@ Appendix A Appendix
(Ramaswami[8]) For $1\geq\beta(2+\alpha\lambda\Delta)$, steady state of $P$ can be found recursively using
@@
Note that, with the random walk presented in $P$, the random walk escapes to infinity when $1\leq\beta(2+\alpha\lambda\Delta)$ which is the same condition as $\beta\geq\frac{1-\beta}{1+(1-\beta)\lambda\Delta}$, i.e., the converse of the ultimate fault tolerance condition. This is not surprising, since $P$ presents the $\Delta$ delay strategy, where the honest chain grows at the slowest possible rate, i.e., $\frac{1}{\Delta+1/\alpha\lambda}$, and is in a race with adversarial chain that has the rate $\beta\lambda$.
In our figures, we use the steady states of $P$ and $\bar{P}$ consistently. We note that the bounds of the two lower bounds visually overlap with each other, hence we refrain from comparing them with plots. As an example, for Bitcoin $6$-block confirmation rule with $\alpha=0.9$, using $P^{\prime}$ results in safety violation lower bound probability of $0.00112412$ whereas $P$ gives $0.00112415$.